OpenStudy (anonymous):

Anyone know how to solve this with TI 84? I have to solve 2e^(-90i) * 6e^(70i). the answer is 12e^(-20i), but my calculator (TI 84 plus silver) says 12e^(-65.9156i). not sure why...

OpenStudy (anonymous):

you should sum 70i and -90i then you get the answer..

OpenStudy (anonymous):

i is located at 2nd, \(\cdot\)

OpenStudy (anonymous):

and yeah, why would one want to do this with a calculator?

OpenStudy (anonymous):

I know how to input the problem, i'm just curious as to why I cannot get the correct form using a calculator.

OpenStudy (anonymous):

i tried with my ti83 ... it just doesnt seem to be well suited to this in any mode.

OpenStudy (anonymous):

is there anyway to convert theta into e^(xi) notation?

OpenStudy (anonymous):

prolly, but i havent played with those identities in soo long that the memory is a bit foggy

OpenStudy (anonymous):

@hazem-wardat

OpenStudy (anonymous):

7.59983044 - 9.28668818 i

OpenStudy (anonymous):

in polar, that is 12e^(-50.704565)

OpenStudy (anonymous):

12e^(-65.9156i) = -11.9799288 - 0.693761651 i

OpenStudy (anonymous):

I know how to get the answer, I just dont know how to get my calculator to show it.

OpenStudy (anonymous):

in polar (12 , -50.70456591